OpenStudy (anonymous):

1. An airline offers a special airfare of $150 to fly from Cincinnati to Washington, DC, a distance of 500 miles. This airline also offered a special airfare of $4000 from Boston to Seattle, a distance of 3000 miles. Write an equation that represents the relationship between the airfare and the flight distance. Then use the equation to find the airfare for a distance of 1200 miles?

OpenStudy (anonymous):

anyone?? Please

OpenStudy (anonymous):

I suggest starting by treating the costs as x values and the distances as y values. Maybe the equation is linear, so you should start by finding a simple ratio between cost and distance

OpenStudy (anonymous):

does not make any sense to me, for first offer is for 150$ a 500 mile trip. if you divide 500/150=3.3333 if you divide 3000/400 it equals 7.5. What do i do if the equations are not linear? help please

OpenStudy (anonymous):

y= (77/50)x- 620, x= distance travelled and y---airfare

OpenStudy (anonymous):

good answer...@akash123

OpenStudy (anonymous):

suppose y= ax +b

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

does it help u?